OpenStudy (anonymous):

if sin a = √5/5 and sin b = √10/10,where a and b are acute find Cos A

OpenStudy (anonymous):

is the problem: if \(\sin(\alpha)=\frac{\sqrt{5}}{5}\) find \(\cos(\alpha)\) ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

can u write the exact question

OpenStudy (anonymous):



OpenStudy (anonymous):

use cosA=sqrt(1-(sina)^2)=sqrt(1-5/25)=sqrt(4/5) =2/sqrt(5)

OpenStudy (anonymous):

there is a picture of an angle whose sine is \(\frac{\sqrt{5}}{5}\) what you need is the adjacent side, which you can find by pythagoras \[\sqrt{5}^2+a^2=5^5\] \[5+a^2=25\] \[a^2=20\] \[a=\sqrt{20}=2\sqrt{5}\]

OpenStudy (anonymous):

the the ratio for cosine would be \[\cos(\alpha)=\frac{2\sqrt{5}}{5}\]